OpenStudy (anonymous):

Calculus Help! A company produces x units of commodity A and y units of commodity B each hour. The company can sell all of its units when commodity A sells for p=25−5x dollars per unit and commodity B sells for q=45−10y dollars per unit. The cost (in dollars) of producing these units is given by the joint-cost function C(x,y)=5xy+3. How much of commodity A and commodity B should be sold in order to maximize profit?

OpenStudy (accessdenied):

Have you made any progress on this problem thus far?

OpenStudy (anonymous):

no

OpenStudy (accessdenied):

The first objective would be finding the profit function to maximize. This is expressed as the revenue from selling subtracting the cost for production. We know our prices for commodity A and B per unit so the revenue can be determined by multiplying units sold (x or y) by its price per unit (p or q). $/unit * unit = $

OpenStudy (anonymous):

So the equation would be (5xy+3)*(25-5x)?

OpenStudy (accessdenied):

Well, we sell x units of commodity A for p=25-5x dollars each. That means the total amount we make is just x*p or x*(25 - 5x). The same occurs for commodity B with y and q. The total revenue, then, is what we make total. x*p + y*q = revenue. Can you see where I get that? Then we know costs, so profit = revenue - costs.

OpenStudy (anonymous):

Ok yeah. so (25x-(5x^2))+(45y-(10y^2))-(5xy+3)

OpenStudy (accessdenied):

Yes! :) So then our goal next is to maximize that function to find optimum values of (x, y). Do you know that process from here?

OpenStudy (anonymous):

Take the partial derivative of x and y?

OpenStudy (accessdenied):

Sounds good to me.

OpenStudy (anonymous):

ok so P(x)=25-10x-5y P(y)=45-20y-5x

OpenStudy (accessdenied):

The partial derivatives are correct to me.

OpenStudy (anonymous):

and then if we want to find a critical point we: P(y)=(45-20y-5x)*-2=(-90+10x+40y) P(x)=(25-10x-5y) Which equals: -65+35y=>y=65/35 ?

OpenStudy (accessdenied):

Yes, that looks good. :) I made a small error myself where I must have wrote q=40-10y instead of 45-10y. But I fixed it and the answers match again!

OpenStudy (anonymous):

Cool! So than I know I have to plug in (65/35) for y but I don't know which "y" to plug it in to lol

OpenStudy (accessdenied):

Both should work to be the same, since it is essentially the intersection point of the two linear equations.

OpenStudy (accessdenied):

45 - 20y - 5x = 0 <-- y=65/35 or 25 - 10x - 5y = 0 <-- y= 65/35 in either case, x will be the same.

OpenStudy (anonymous):

Alright so 25-10x-5(65/35)=approx 1.571428517!!!

OpenStudy (accessdenied):

Looks good to me! Could also be represented in fractions for the nicer numbers: x=11/7, y=13/7. I am thinking that is the answer, although I don't know if the problem wants to assume x and y have to be integers (like, you can't sell partial commodity). I might be overcomplicating that. lol

OpenStudy (anonymous):

Yes it is the right answer (the homework is online so we know instantly if it's right)!

OpenStudy (accessdenied):

Alright, great! :P

OpenStudy (anonymous):

Thank you, So one final quick question... Would I do about the same thing for: "A manufacturer is planning to sell a new product at the price of $290 per unit and estimates that if x thousand dollars is spent on development and y thousand dollars is spent on promotion, consumers will buy approximately (250y/y+4)+(500x/x+3) units of the product. If manufacturing costs $120 per unit, how much should the manufacturer spend on development and how much on promotion to generate the largest possible profit from the sale of this product?"

OpenStudy (accessdenied):

Your goal of the equation "profit = revenue - cost" is the same, it just looks like you have to find revenue and cost differently from the first one.

OpenStudy (anonymous):

So, P would=1700((25y)/y+4)+((50x)/x+3)-1000x-1000y

OpenStudy (accessdenied):

Is that written like: \[ 1700 \frac{25y}{y + 4} \cdots \] I think it looks correct from that.

OpenStudy (anonymous):

Ok so then the partial derivative would be?

OpenStudy (accessdenied):

There is a quotient of two polynomials, so you would use quotient rule for derivatives to find that part.

OpenStudy (anonymous):

would P(x)=-(1000 (-246+6 x+x^2))/(3+x)^2 would P(y)=-(1000 (-154+8 y+y^2))/(4+y)^2?

OpenStudy (accessdenied):

Sorry, I wasn't getting that for an answer. I need to retry my steps...

OpenStudy (anonymous):

Ok no problem, If you can't get it it's no problem because my homework was due at 11 lol. Thanks for all the help though!

OpenStudy (accessdenied):

I think you are correct, I just seem to be getting a weird answer. Might be the sleepiness. lol

OpenStudy (anonymous):

Alright I'll let you go. Again, thanks for the help!

OpenStudy (accessdenied):

No problem! I think once you got the function for profit, you are fine! :D

OpenStudy (accessdenied):

I glanced over my work and found I had written 50 instead of 150. Fixing that, I got the same partial derivatives you had. lol